OpenStudy (vickyisthesmartone123):

The altitude of a plane as it begins to descend is shown by the linear equation a = 30,000 – 1500m, where a is the plane’s altitude, in feet, and m is minutes. What is the plane’s altitude when m = 12?

OpenStudy (vickyisthesmartone123):

@KyanTheDoodle

OpenStudy (vickyisthesmartone123):

@Brookie105

OpenStudy (vickyisthesmartone123):

Help please pronto i need to turn in this test in 5 minutes!! anyone plz???.......

OpenStudy (superdavesuper):

put m=12 into the eqn n calculate a...

OpenStudy (vickyisthesmartone123):

can u tell me the answer i aint got the time boy...plz i dont want to sound dramamtic but..plz?

OpenStudy (superdavesuper):

sorry no direct ans on OS here

OpenStudy (vickyisthesmartone123):

ok fine..thyanks i will jsut solve it on my own bye!!

OpenStudy (vickyisthesmartone123):

@KyanTheDoodle

OpenStudy (vickyisthesmartone123):

whick way do i subtract it <---or----->way?

TheSmartOne (thesmartone):

You will get your answer faster if you cooperate instead of asking for a direct answer or finding someone who will give you a direct answer.

OpenStudy (vickyisthesmartone123):

ok fine i just need to know if you subtrzact 1512 by 30000 or the other way

OpenStudy (vickyisthesmartone123):

or 30000-1512

TheSmartOne (thesmartone):

a = 30,000 – 1500m Plug in m = 12 a = 30,000 - 1,500(12) First you have to multiply. What is 1500 x 12 = ?

OpenStudy (vickyisthesmartone123):

18000

TheSmartOne (thesmartone):

And 30,000 - 18,000 = ?

OpenStudy (vickyisthesmartone123):

12000 so thats my answer

TheSmartOne (thesmartone):

yup! :D

OpenStudy (vickyisthesmartone123):

ok thanks bye!:)

TheSmartOne (thesmartone):

umm bye

OpenStudy (vickyisthesmartone123):

wait i need help i tagged u to a question